# Continuous Urban Change Detection from Satellite Image Time Series with Temporal Feature Refinement and Multi-Task Integration

### v2

@@ Abstract
Urbanization advances at unprecedented rates, resulting in negative effects on the environment and human well-being. Remote sensing has the potential to mitigate these effects by supporting sustainable development strategies with accurate information on urban growth. Deep learning-based methods have achieved promising urban change detection results from optical satellite image pairs using convolutional neural networks (ConvNets), transformers, and a multi-task learning setup. However, transformers have not been leveraged for urban change detection with multi-temporal data, i.e., $>2$ images, and multi-task learning methods lack integration approaches that combine change and segmentation outputs. To fill this research gap, we propose a continuous urban change detection method that identifies changes in each consecutive image pair of a satellite image time series (SITS). Specifically, we propose a temporal feature refinement (TFR) module that utilizes self-attention to improve ConvNet-based multi-temporal building representations. Furthermore, we propose a multi-task integration (MTI) module that utilizes Markov networks to find an optimal building map time series based on segmentation and dense change outputs. The proposed method effectively identifies urban changes based on high-resolution SITS acquired by the PlanetScope constellation (F1 score 0.551) and Gaofen-2 (F1 score 0.440). Moreover, our experiments on two challenging datasets demonstrate the effectiveness of the proposed method compared to bi-temporal and multi-temporal urban change detection and segmentation methods. Code is available on GitHub: https://github.com/SebastianHafner/ContUrbanCD.
@@ I Introduction
In recent years, deep learning has been continuously replacing traditional change detection methods [8, 9, 10, 11]. Specifically, deep convolutional neural networks (ConvNets) have been used extensively for change detection in bi-temporal optical satellite image pairs (see Fig. 1(a)). The simplest way of adapting common ConvNets such as U-Net [12] for change detection is with an input-level fusion (or early fusion [13]) strategy, referring to the concatenation of image pairs before passing them to a ConvNet. Contrarily, late fusion strategies typically process images separately in a Siamese network consisting of two ConvNets with shared weights. Extracted bi-temporal features are then fused using concatenation or absolute differencing [13, 14]. Since Siamese networks are generally considered preferable to input-level fusion strategies, multiple studies developed modules that are incorporated into Siamese network architectures to improve feature representations [15, 16, 17]. For example, Chen et al. [18, 19] proposed to refine features extracted by ConvNets from very high-resolution (VHR) imagery using a transformer-based module, alleviating the limited long-range context modeling capability of convolutions with self-attention. Since then, self-attention has become a popular mechanism for capturing long-range contextual information in VHR change detection [20, 17, 21, 22].
@@
In recent years, high-resolution (i.e., 1 – 10 m) satellite image time series (SITS) have become increasingly available [23]. Those data have enabled a shift from detecting land cover changes in image pairs acquired years apart to continuous annual and sub-annual change detection [24]. In contrast, urban change detection methods are predominantly designed for bi-temporal change detection from image pairs acquired multiple years apart [25]. However, considering the unprecedented rate of global urbanization [26], it is essential to develop a new suite of methods that detect urban changes continuously. While continuous change detection can be achieved by applying a bi-temporal model to consecutive image pairs in SITS (see Fig. 1(c)), this approach fails to exploit multi-temporal, i.e., $>$ 2 images, information. Furthermore, recent studies [27, 28] have demonstrated the effectiveness of multi-temporal change detection models which predict changes between the first and last images of a SITS (see Fig. 1(b)). For example, multi-temporal information can help to reduce commission errors from registration errors, illumination differences, or other types of change unrelated to the problem of interest [28]. In addition, it can mitigate the effect of cloud artifacts in single images [27]. Existing multi-temporal change detection methods employ either recurrent layers, such as long short-term memory (LSTM) layers [29], or 3D convolutional layers to model temporal information [27, 28, 30, 31]. While these layers effectively model short-range temporal dependencies in time series data, the self-attention mechanism can explicitly model temporal dependencies across all timestamps of a time series [32]. Thus, several recent segmentation methods for SITS employ the self-attention mechanism to explicitly model temporal dependencies across all timestamps of a time series [33, 34, 35]. However, the temporal models in these methods collapse the temporal dimension, resulting in a single output feature. Therefore, they do not facilitate continuous urban change detection which requires the full temporal information to produce change maps between each consecutive image pair in the SITS.
@@
In this paper, we propose a continuous urban change detection method (see Fig. 1(d)) and explore two research gaps in the current literature, namely (1) the use of the self-attention mechanism to model multi-temporal information in urban change detection and (2) the integration of segmentation and change predictions in multi-task learning setups. Specifically, we propose a new network architecture that relies on convolutions to extract multi-temporal building representations and employs self-attention to model temporal dependencies in feature space while preserving the temporal dimension. We also propose a novel integration approach that determines the optimal building segmentation for each image in a time series based on the multi-task network outputs. The effectiveness of the proposed architecture and integration approach is demonstrated on two urban change detection datasets featuring high-resolution optical SITS, namely the SpaceNet 7 dataset [41] and the Wuhan Urban Semantic Understanding (WUSU) dataset [42].
@@
  • Experiments on two datasets, namely SpaceNet 7 and WUSU, show that the proposed continuous urban change detection method is more effective than related methods.
@@ II-A Bi-temporal change detection
In recent years, a plethora of deep learning-based bi-temporal change detection methods have been proposed. Most of these works focus on developing new Siamese network architectures and/or training strategies. Initially, Daudt et al. [13, 14] proposed two Siamese ConvNet architectures for change detection. The Siam-Diff and Siam-Conc architectures employ encoders with shared weights for feature extraction from bi-temporal high-resolution image pairs and combine the corresponding feature maps using a subtraction and concatenation strategy, respectively. While encoders and decoders in these models follow the U-Net architecture [12], Fang et al. [15] incorporated a nested U-Net (i.e., UNet++ [43]) into a Siamese network to maintain high-resolution fine-grained representations through dense skip connections. Many works also improved Siamese networks by incorporating different modules into the architecture. For example, an ensemble channel attention module was proposed for feature refinement in [15], and a new spatial pyramid pooling block was utilized in [16] to preserve shapes of change areas.
@@ III-A Overview
As illustrated in Fig. 2, the main components of the proposed method are a ConvNet-based encoder, transformer-based temporal feature refinement (TFR) modules, a change feature (CF) module, two task-specific ConvNet-based decoders, and a Markovian multi-task (MTI) integration module. The following summarizes the urbanization monitoring process of the proposed method for a time series of satellite images:
@@ III-B Satellite image time series encoding
where $e(.)$ represents the encoder, and subscript ${\rm seg}$ indicates that the feature maps contain representations for building segmentation.
@@ III-D Change feature extraction
We refer to the scenario when all possible combinations of image pairs are considered as dense. However, we also investigate sparser settings such as adjacent, considering only changes between temporally adjacent images ($N_{\rm adjacent}=T-1$), and cyclic, adding the changes between the first and last image to the adjacent setting ($N_{\rm cyclic}=T$) (see Fig. 5(a)5(c) for visualizations of these settings in Markov networks with $T=5$). We provide ablation results on the different edge settings in Tables III and V. Finally, it should be noted that the CF module does not have any trainable parameters.
@@ III-E Multi-task decoding
Two separate decoders are deployed to convert the temporally enriched segmentation feature maps and the change feature maps to building outputs and urban change outputs, respectively. Formally, we obtain $T$ built-up area segmentation outputs $\mathbf{O}_{\rm seg}\in(0,1)^{T\times H\times W}$ with the segmentation decoder $d_{\rm seg}(.)$ as follows:
@@
Furthermore, we obtain $N$ change detection outputs $\mathbf{O}_{\rm ch}\in(0,1)^{N\times H\times W}$ with the change decoder $d_{\rm ch}(.)$ as follows:
@@ III-G Multi-task integration
To define a global model from the local interactions defined in the parameterization of the Markov network, we take the product of the local factors and then normalize it. Once the distribution is defined, we perform a maximum a posteriori query to find the optimal state assignment for each node in the graph. To that end, inference is performed using the belief propagation algorithm (see Algorithm 10.4 in [49]). Due to the absence of loops in the graph, belief propagation provides an exact solution. Finally, it should be noted that the MTI module does not contain any trainable parameters and is only deployed during inference.
@@ IV-A Datasets
WUSU: The Wuhan Urban Semantic Understanding (WUSU) dataset features tri-temporal high-resolution Gaofen-2 images covering Wuhan’s Hongshan District and Jiang’an District [53]. The preprocessing workflow of the satellite images includes orthographic correction and multi-temporal image registration. Furthermore, the four multi-spectral bands acquired at 4 m spatial resolution are pansharpened to a spatial resolution of 1 m, resulting in images of size 6358 x 6382 and 7025 x 5500 pixels for Hongshan District and Jiang’an District, respectively. In addition to the Gaofen-2 images, the WUSU dataset provides corresponding land-use/land-cover (LULC) labels, including manually refined building annotations (Class 2 Low building and Class 3 High building). Since the proposed method requires binary building labels, Class 2 and Class 3 were remapped to the foreground class, whereas all other classes were remapped to the background class. We follow the within-scene split recommended by the authors, using the top halves of the six images for the test set and the bottom halves for the training set that was further divided into training (90 %) and validation (10 %) tiles.
@@ IV-B Baseline and benchmark methods
We selected a comprehensive set of baseline and benchmark methods for quantitative and qualitative comparisons with the proposed methods. These selected methods grouped into two categories are listed below.
@@ V-A SpaceNet 7
The quantitative results for the SpaceNet 7 dataset are listed in Table I. The proposed method achieved the highest F1 scores and IoU values for both urban change detection tasks (i.e., bi-temporal and continuous). Several multi-temporal models (L-UNet, Multi-Task L-UNet, and U-TempoNet) outperform bi-temporal change detection methods on the bi-temporal task while others are less effective (U-TAE and TSViT). Among the bi-temporal methods, ScartchFormer and the ConvNet-based methods Siam-Diff and SNUNet achieved the highest accuracy values. For building segmentation, the proposed method also outperformed the other multi-task methods including DTCDSCN and Multi-Task L-UNet.
TABLE I: Quantitative results on the SpaceNet 7 test areas. The best and second-best performances are highlighted in red and blue, respectively. ”-” denotes that the accuracy metric does not apply to a specific method since the corresponding variable is not predicted.
@@ V-B WUSU
The quantitative results for the WUSU datasets are listed in Table II. In general, change detection performance on the WUSU dataset is lower than on the SpaceNet 7 dataset. For bi-temporal change detection, only ScratchFormer, L-UNet, and Multi-Task L-UNet exceed an F1 score of 0.275 and an IoU value of 0.160. In comparison, accuracy values for the continuous change detection task are slightly higher, except for the proposed method. Overall, our method outperforms all other methods on both change detection tasks. This also applied to the building segmentation task, where the proposed method achieved an F1 score of 0.663 and an IoU value of 0.496.
TABLE II: Quantitative results on the WUSU test areas. The best and second-best performances are highlighted in red and blue, respectively. ”-” denotes that the accuracy metric does not apply to a specific method since the corresponding variable is not predicted.
@@
Fig. 8 and 9 show qualitative change detection and building segmentation results for two WUSU test sites located in Wuhan’s Jiang’an District in China. The Gaofen-2 SITS for both sites highlight the challenging nature of the WUSU dataset. In particular, the images were acquired under different atmospheric conditions and they have large off-nadir angles and contain shadows. Despite that, the change detection results produced by the proposed method show good agreement with the label, especially in comparison with the competing methods. As for SpaceNet 7, the proposed method achieved a high level of consistency between the continuous change detection outputs (rows two and three) and the change output between the first and last image (top row). In contrast, the bi-temporal change detection methods identified changes between the first and last images (ch $t1\rightleftharpoons t3$) that are present in neither of the continuous change rows (ch $t1\rightleftharpoons t2$ and ch $t2\rightleftharpoons t3$). Finally, the bottom row demonstrates that all methods accurately map buildings, but the proposed method achieved more detailed building delineations than its competitors. It should also be noted that the label does not distinguish individual buildings in very dense built-up areas (e.g., the bottom left area in Fig. 8).
@@
V-C Ablation study
Loss function: Table III shows the ablation results for different change loss edge settings (first-last, adjacent, cyclic, and dense) and the segmentation loss. It should be noted that the settings adjacent, cyclic, and dense all require building annotations for each image in the time series as labels (see Section III-F, whereas first-last only requires building annotations for the first and last images. In general, considering additional change edges in the loss function improves continuous change detection performance. However, a single change loss term suffices for the bi-temporal change detection task. The segmentation loss term improves performance for both change detection tasks, which holds for all change loss settings. Therefore, the optimal loss setting consists of a change loss with dense edges combined with a segmentation loss.
TABLE III: Ablation results for the loss function with different edge settings for the change loss term. The best and second-best performances are highlighted in red and blue, respectively.

### v3

@@ Abstract
Urbanization advances at unprecedented rates, leading to negative environmental and societal impacts. Remote sensing can help mitigate these effects by supporting sustainable development strategies with accurate information on urban growth. Deep learning-based methods have achieved promising urban change detection results from optical satellite image pairs using convolutional neural networks (ConvNets), transformers, and a multi-task learning setup. However, bi-temporal methods are limited for continuous urban change detection, i.e., the detection of changes in consecutive image pairs of satellite image time series (SITS), as they fail to fully exploit multi-temporal data ($>$ 2 images). Existing multi-temporal change detection methods, on the other hand, collapse the temporal dimension, restricting their ability to capture continuous urban changes. Additionally, multi-task learning methods lack integration approaches that combine change and segmentation outputs. To address these challenges, we propose a continuous urban change detection framework incorporating two key modules. The temporal feature refinement (TFR) module employs self-attention to improve ConvNet-based multi-temporal building representations. The temporal dimension is preserved in the TFR module, enabling the detection of continuous changes. The multi-task integration (MTI) module utilizes Markov networks to find an optimal building map time series based on segmentation and dense change outputs. The proposed framework effectively identifies urban changes based on high-resolution SITS acquired by the PlanetScope constellation (F1 score 0.551), Gaofen-2 (F1 score 0.440), and WorldView-2 (F1 score 0.543). Moreover, our experiments on three challenging datasets demonstrate the effectiveness of the proposed framework compared to bi-temporal and multi-temporal urban change detection and segmentation methods. Code is available on GitHub: https://github.com/SebastianHafner/ContUrbanCD.
@@ I Introduction
In recent years, deep learning has been continuously replacing traditional change detection methods [8, 9, 10, 11]. Specifically, deep convolutional neural networks (ConvNets) have been used extensively for change detection in bi-temporal optical satellite image pairs (see Fig. 1(a)). The simplest way of adapting common ConvNets such as U-Net [12] for change detection is with an input-level fusion (or early fusion [13]) strategy, referring to the concatenation of image pairs before passing them to a ConvNet. Contrarily, late fusion strategies typically process images separately in a Siamese network consisting of two ConvNets with shared weights. Extracted bi-temporal features are then fused using concatenation or absolute differencing [13, 14]. Since Siamese networks are generally considered preferable to input-level fusion strategies, multiple studies developed modules that are incorporated into Siamese network architectures to improve feature representations [15, 16, 17]. For example, Chen et al. [18, 19] proposed to refine features extracted by ConvNets from very high-resolution (VHR) imagery using a transformer-based module, alleviating the limited long-range context modeling capability of convolutions with self-attention. Since then, self-attention has become a popular mechanism for capturing long-range spatial dependencies in VHR change detection [20, 17, 21, 22].
@@
In recent years, high-resolution (i.e., 1 – 10 m) satellite image time series (SITS) have become increasingly available [23]. Those data have enabled a shift from detecting land cover changes in image pairs acquired years apart to continuous annual and sub-annual change detection [24]. In contrast, urban change detection methods are predominantly designed for bi-temporal change detection from image pairs acquired multiple years apart [25]. However, considering the unprecedented rate of global urbanization [26], it is essential to develop a new suite of methods that detect urban changes continuously. While continuous change detection can be achieved by applying a bi-temporal model to consecutive image pairs in SITS (see Fig. 1(c)), this approach fails to exploit multi-temporal, i.e., $>$ 2 images, information. Furthermore, recent studies [27, 28] have demonstrated the effectiveness of multi-temporal change detection models that predict changes between the first and last images of a SITS (see Fig. 1(b)). For example, multi-temporal information can help to reduce commission errors from registration errors, illumination differences, or other types of change unrelated to the problem of interest [28]. In addition, it can mitigate the effect of cloud artifacts in single images [27]. Existing multi-temporal change detection methods employ either recurrent layers, such as long short-term memory (LSTM) layers [29], or 3D convolutional layers to model temporal information [27, 28, 30, 31]. While these layers effectively model short-range temporal dependencies in time series data, the self-attention mechanism can explicitly model temporal dependencies across all timestamps of a time series [32]. Thus, several recent segmentation methods for SITS employ the self-attention mechanism to explicitly model temporal dependencies across all timestamps of a time series [33, 34]. However, the temporal models in these methods collapse the temporal dimension, resulting in a single output feature. Therefore, they do not facilitate continuous urban change detection, which requires the full temporal information to produce change maps between each consecutive image pair in the SITS.
@@
In this paper, we propose a continuous urban change detection method (see Fig. 1(d)) and explore two research gaps in the current literature, namely (1) the modeling of multi-temporal information using self-attention for continuous urban change detection and (2) the integration of segmentation and change predictions in multi-task learning setups. Specifically, we propose a new network architecture that relies on convolutions to extract multi-temporal building representations and employs self-attention to model temporal dependencies in feature space while preserving the temporal dimension. We also propose a novel integration approach that determines the optimal building segmentation for each image in a time series based on the multi-task network outputs. The effectiveness of the proposed architecture and integration approach is demonstrated on three urban change detection datasets featuring high-resolution optical SITS.
@@
  • Experiments on three datasets, namely SpaceNet 7, WUSU, and TSCD, show that the proposed continuous urban change detection method is more effective than related methods.
@@ II-A Bi-temporal change detection
In recent years, a plethora of deep learning-based bi-temporal change detection methods have been proposed. Most of these works focus on developing new Siamese network architectures and/or training strategies. Initially, Daudt et al. [13, 14] proposed two Siamese ConvNet architectures for change detection. The Siam-Diff and Siam-Conc architectures employ encoders with shared weights for feature extraction from bi-temporal high-resolution image pairs and combine the corresponding feature maps using a subtraction and concatenation strategy, respectively. While encoders and decoders in these models follow the U-Net architecture [12], Fang et al. [15] incorporated a nested U-Net (i.e., UNet++ [40]) into a Siamese network to maintain high-resolution, fine-grained representations through dense skip connections. Many works also improved Siamese networks by incorporating different modules into the architecture. For example, an ensemble channel attention module was proposed for feature refinement in [15], and a new spatial pyramid pooling block was utilized in [16] to preserve shapes of change areas.
@@ III-A Overview
As illustrated in Fig. 2, the main components of the proposed method are a ConvNet-based encoder, transformer-based temporal feature refinement (TFR) modules, a change feature (CF) module, two task-specific ConvNet-based decoders, and a Markovian multi-task integration (MTI) module. The following summarizes the urbanization monitoring process of the proposed method for a time series of satellite images:
@@ III-B Satellite image time series encoding
where $e(\cdot)$ represents the encoder, and subscript ${\rm seg}$ indicates that the feature maps contain representations for building segmentation.
@@ III-D Change feature extraction
We refer to the scenario when all possible combinations of image pairs are considered as dense. However, we also investigate sparser settings such as adjacent, considering only changes between temporally adjacent images ($N_{\rm adjacent}=T-1$), and cyclic, adding the changes between the first and last image to the adjacent setting ($N_{\rm cyclic}=T$) (see Fig. 5(a)5(c) for visualizations of these settings in Markov networks with $T=5$). We provide ablation results on the different edge settings in Tables V and VII. Finally, it should be noted that the CF module does not have any trainable parameters.
@@ III-E Multi-task decoding
Two separate decoders are deployed to convert the temporally enriched segmentation feature maps and the change feature maps to building outputs and urban change outputs, respectively. Formally, we obtain $T$ built-up area segmentation outputs $\mathbf{O}_{\rm seg}\in(0,1)^{T\times H\times W}$ with the segmentation decoder $d_{\rm seg}(\cdot)$ as follows:
@@
Furthermore, we obtain $N$ change detection outputs $\mathbf{O}_{\rm ch}\in(0,1)^{N\times H\times W}$ with the change decoder $d_{\rm ch}(\cdot)$ as follows:
@@ III-G Multi-task integration
To define a global model from the local interactions defined in the parameterization of the Markov network, we take the product of the local factors and then normalize it. Once the distribution is defined, we perform a maximum a posteriori query to find the optimal state assignment for each node in the graph. The optimal state assignment corresponds to the configuration that minimizes the overall energy determined by the node and edge potentials assigned to the graph, as defined in Equations 10 and 11. Therefore, the resulting state assignment for the nodes is optimal with respect to the potentials obtained from the network outputs but not necessarily with respect to the segmentation and change labels. We perform inference using the belief propagation algorithm (see Algorithm 10.4 in [47]). Due to the absence of loops in the graph, belief propagation provides an exact solution. Finally, it should be noted that the MTI module does not contain any trainable parameters and is only deployed during inference.
@@ IV-A Datasets
A diverse set of multi-temporal datasets is used to evaluate the proposed method. Table I compares key characteristics of these datasets, and detailed descriptions are provided in the following paragraphs.
TABLE I: Overview of dataset characteristics for SpaceNet 7, the Wuhan Urban Semantic Understanding (WUSU) dataset, and the Time Series Change Detection (TSCD) dataset.
Dataset Location Satellite Resolution
Spectral Spatial Temporal
SpaceNet 7 Global (60 sites) PlanetScope 3 bands (RGB) 4 m 5 images* (2017 to 2020)
WUSU Wuhan, China Gaofen-2 4 bands (RGB + NIR) 1 m 3 images (2015, 2016, 2018)
TSCD Chengdu, China WorldView-2 3 bands (RGB) 0.5 m 4 images (2016, 2018, 2020, 2022)
    Approximately 24 monthly timestamps acquired between 2017 and 2020 are available for each site in the dataset.
@@
WUSU: The Wuhan Urban Semantic Understanding (WUSU) dataset features tri-temporal high-resolution Gaofen-2 images covering two districts in Wuhan (Hubei Province, central China) in 2015, 2016, and 2018 [52]. The preprocessing workflow of the satellite images includes orthographic correction and multi-temporal image registration. Furthermore, the four multi-spectral bands acquired at 4 m spatial resolution are pansharpened to a spatial resolution of 1 m, resulting in images of size 6358 x 6382 and 7025 x 5500 pixels for Hongshan District and Jiang’an District, respectively. In addition to the Gaofen-2 images, the WUSU dataset provides corresponding land-use/land-cover (LULC) labels, including manually refined building annotations (Class 2 Low building and Class 3 High building). Since the proposed method requires binary building labels, Class 2 and Class 3 were remapped to the foreground class, whereas all other classes were remapped to the background class. We follow the within-scene split recommended by the authors, using the top halves of the six images for the test set and the bottom halves for the training set that was further divided into training (90 %) and validation (10 %) tiles.
TSCD: The Time Series Change Detection (TSCD) dataset features bi-annual WorldView-2 satellite images acquired over Chengdu (Sichuan Province) between 2016 and 2022 [53]. The images have a resolution of approximately 0.5 m and are split into 512 x 512 pixel tiles. The tiles are divided into a training, validation, and test set. The TSCD dataset provides building change labels for each adjacent image pair (2016–2018, 2018–2020, 2020–2022). We derived change labels for an arbitrary image pair from the time series by considering all adjacent change labels connecting this pair and computing the number of changes. An odd number of adjacent changes indicates change between the image pair, whereas an even number indicates no change.
@@ IV-B Baseline and benchmark methods
We selected a comprehensive set of baseline and benchmark methods for quantitative and qualitative comparisons with the proposed methods. These selected methods are grouped into two categories, which are listed below.
@@ V-A SpaceNet 7
The quantitative results for the SpaceNet 7 dataset are listed in Table II. The proposed method achieved the highest F1 scores and IoU values for both urban change detection tasks (i.e., bi-temporal and continuous). Several multi-temporal models (L-UNet, Multi-Task L-UNet, and U-TempoNet) outperform bi-temporal change detection methods on the bi-temporal task while others are less effective (U-TAE and TSViT). Among the bi-temporal methods, ScartchFormer and the ConvNet-based methods Siam-Diff and SNUNet achieved the highest accuracy values. For building segmentation, the proposed method also outperformed the other multi-task methods including DTCDSCN and Multi-Task L-UNet.
TABLE II: Quantitative results on the SpaceNet 7 test areas. The best and second-best performances are highlighted in red and blue, respectively. ”-” denotes that the accuracy metric does not apply to a specific method since the corresponding variable is not predicted.
@@ V-B WUSU
Change detection performance on the WUSU dataset is lower than on the SpaceNet 7 dataset (Table III). For bi-temporal change detection, only ScratchFormer, L-UNet, and Multi-Task L-UNet exceed an F1 score of 0.275 and an IoU value of 0.160. In comparison, accuracy values for the continuous change detection task are slightly higher, except for the proposed method. Overall, our method outperformed all other methods on both change detection tasks. This also applied to the building segmentation task, where the proposed method achieved an F1 score of 0.663 and an IoU value of 0.496.
TABLE III: Quantitative results on the WUSU test areas. The best and second-best performances are highlighted in red and blue, respectively. ”-” denotes that the accuracy metric does not apply to a specific method since the corresponding variable is not predicted.
@@
Fig. 8 and 9 show qualitative change detection and building segmentation results for two WUSU test sites located in Wuhan’s Jiang’an District in China. The change detection results produced by the proposed method show good agreement with the label, especially in comparison with the competing methods. As for SpaceNet 7, the proposed method achieved a high level of consistency between the continuous change detection outputs (rows two and three) and the change output between the first and last image (top row). In contrast, the bi-temporal change detection methods identified changes between the first and last images (ch $t1\rightleftharpoons t3$) that are present in neither of the continuous change rows (ch $t1\rightleftharpoons t2$ and ch $t2\rightleftharpoons t3$). Finally, the bottom row demonstrates that all methods accurately map buildings, but the proposed method achieved more detailed building delineations than its competitors. It should also be noted that the label does not distinguish individual buildings in very dense built-up areas (e.g., the bottom left area in Fig. 8).
@@
V-C TSCD
The quantitative results for the TSCD dataset are listed in Table IV. Multi-temporal methods outperformed bi-temporal methods on detecting changes between the first and last timestamps of SITS. On the continuous task, bi-temporal methods capable of modeling long-range spatial dependencies in VHR imagery (BIT, AMTNet, and ScratchFormer) outperformed Siam-Diff and SNUNet. However, our method achieved the highest performance on both tasks despite not leveraging spatial attention.
TABLE IV: Quantitative results on the TSCD test areas. The best and second-best performances are highlighted in red and blue, respectively. ”-” denotes that the accuracy metric does not apply to a specific method since the corresponding variable is not predicted.
Method Change detection
Bi-temporal Continuous
(F1 / IoU / OA) (F1 / IoU / OA)
Siam-Diff 0.211 / 0.118 / 89.3 0.284 / 0.166 / 93.2
SNUNet 0.220 / 0.123 / 89.4 0.328 / 0.197 / 92.6
BIT 0.203 / 0.114 / 90.0 0.363 / 0.222 / 94.5
AMTNet 0.298 / 0.178 / 91.1 0.398 / 0.249 / 94.9
ScratchFormer 0.270 / 0.157 / 91.7 0.504 / 0.340 / 95.9
L-UNet 0.401 / 0.251 / 84.6 -
U-TAE 0.415 / 0.262 / 81.6 -
TSViT 0.355 / 0.216 / 79.3 -
U-TempoNet 0.435 / 0.279 / 85.1 -
Proposed 0.543 / 0.373 / 93.7 0.573 / 0.402 / 96.6
V-D Ablation study
Loss function: Table V shows the ablation results for different change loss edge settings (first-last, adjacent, cyclic, and dense) and the segmentation loss. Multi-task integration was disabled for this experiment to isolate the effect of the loss function on network performance. It should also be noted that the settings adjacent, cyclic, and dense all require building annotations for each image in the time series as labels (see Section III-F), whereas first-last only requires building annotations for the first and last images. Considering additional change edges in the loss function generally improves continuous change detection performance. In contrast, a single change loss term suffices for the bi-temporal change detection task. The segmentation loss term improves performance for both change detection tasks, which holds for all change loss settings. Therefore, the optimal loss setting consists of a change loss with dense edges combined with a segmentation loss.
TABLE V: Ablation results for the loss function with different edge settings for the change loss term. The best and second-best performances on the SpaceNet 7 (SN7), WUSU, and TSCD datasets are highlighted in red and blue, respectively.